# LaMET’s Asymptotic Extrapolation vs. Inverse Problem

### v1

@@ IV.4 Asymptotic Analysis Provides Realistic Error Estimate Regardless of Data Quality
The lattice correlation functions include statistical errors that are correlated and will propagate in the FT. Since FT is a linear operation, there are mature statistical methods to handle the error propagation [220]. In addition to statistical errors, the FT also includes model uncertainty arising from the extrapolation. Another main point advocated in Ref. [1] is that when the FT uncertainty from extrapolation is non-negligible, it cannot be reliably quantified. We disagree. As we show here, the uncertainty can be estimated within asymptotic extrapolation and does not constitute the IP encountered in SDF.
As explained above, a non-negligible model uncertainty is the outcome of insufficient data quality in the asymptotic region, which can be reduced by improving the data precision. To quantify this uncertainty, it suffices to show that it is both bounded from above and systematically reducible. The upper bound can be estimated by evaluating the contribution of the extrapolated region to the Fourier integral. One bound has been derived in the appendix of Ref. [65] as
@@
where $|h(z,P;\lambda_{L})|_{\rm max}$ is the maximal value of $|h(z,P)|$ for $\lambda_{L}\leq\lambda<\infty$  [^3] We have slightly adapted the derivation in Ref. [65], which assumed that $|h(\lambda,P)|$ decays monotonically for $\lambda>\lambda_{L}$. In our version, the derivation also allows for oscillations in this region., and $N_{x}$ is an integer such that, after sufficiently long decay, the contribution to the Fourier integral from $\lambda>\lambda_{L}+N_{x}\cdot(2\pi)/x$ becomes negligible.
@@
The upper bound in Eq. (9) applies to both exponential and algebraic decays, typically requiring $N_{x}\sim\mathcal{O}(1)$ in the former and $N_{x}\gg 1$ in the latter to saturate the Fourier integral. This contrast is expected: for the same $\lambda_{L}$, the extrapolated region contributes much less to the Fourier integral in the case of exponential decay, leading to a correspondingly smaller model uncertainty. For exponential decays, this upper bound is usually comparable to or smaller than the other uncertainties in the LaMET analyses, making it a reasonable estimate of the FT model uncertainty. Moreover, it excessively estimates the true uncertainty, as it assumes that $|h(z,P)|$ can vary arbitrarily between 0 and its maximum value $|h(z,P;\lambda_{L})|_{\rm max}$ at each $z$. With a more accurate derivation of the asymptotic behavior of spatial correlations [208], a significantly tighter bound on the error budget can be established.
@@
Since the “Exact exp” and RBF kernels in Ref. [1] all included the exponential decay, it is appropriate to apply the upper bound in Eq. (9). The FT model uncertainty can be estimated by the variation of central values across different models, which should persist even when the statistical errors approach zero. Because the extrapolations using the RBF kernel do not satisfy the third physical condition in Sec. III.3, we set $|h(z,P;\lambda_{L})|_{\rm max}=0.1$, reflecting the amplitude of the matrix elements in the range $0.75\,\text{fm}<z<1.03\,\text{fm}$. At $x=0.5$, taking $N_{x}=1$ is sufficient to saturate the Fourier integral, yielding an upper bound of $\sim 0.25$ on the model uncertainty. Meanwhile, according to Figure 3 of Ref. [1], the quasi-PDFs obtained from different models are consistent within errors at finite $x$, despite that their central values differ by an amount that generally decreases with increasing $x$, consistent with the behavior of the upper bound in Eq. (9). See Figure 5 of Ref. [1]. At $x=0.5$, the spread among central values is around 0.2, implying an uncertainty of about 0.1, which is still bounded by the estimate from Eq. (9).
@@
Moreover, with the precise data in Fig. 4(a), the quasi-PDFs from different extrapolations show significantly better agreement, as illustrated in Figs. 4(b) and 4(c). The variation around the mean of the Exact exp model is less than 0.05 at $x=0.5$, which is again bounded by the estimated uncertainty of $\sim 0.06$, obtained using $|h(z,P;\lambda_{L})|_{\rm max}=0.025$ and $N_{x}=1$. Notably, the deviations of GPR fits remain outside the error bands of the Exact exp extrapolation, indicating that they are not negligible. However, this reflects a problem of the GPR methods themselves, as they do not meet the physical conditions for asymptotic extrapolation.
@@ IV.5 Exponential Decay is the Key to Controlling FT Errors in LaMET
Ref. [1] also made the claim that exponential decay plays a minimal role in controlling the FT at moderate to large $x$, so that the inverse problem in LaMET and SDF is essentially the same. This claim is based on the observation that different $z$- and $x$-space reconstruction methods yield similar results for the quasi-PDF, regardless of whether exponential decay is imposed.
However, the logic behind that conclusion is completely reversed. It is the asymptotic exponential decay of the pseudo data—not the extrapolation model—that ensures control of the FT. This is why similar results can still be obtained even when the decay behavior is not explicitly built into the model, exactly the first issue with the data-driven methods discussed in Sec. IV.3. To match the exponentially decaying data, the extrapolation must also satisfy the corresponding physical constraints, otherwise the systematic errors cannot be reliably estimated.
In contrast, the matrix elements in SDF are limited to $\lambda\lesssim 4.0$ and decay much more slowly, which inevitably leads to significantly larger variations across different extrapolation models, as shown in Fig. 2. Imposing any model, even those incorporating exponential decay, does not resolve the IP in this case.
@@ V Conclusion
In this comment, we have explained why the asymptotic extrapolation in LaMET enables a reliable extrapolation of the long-range lattice matrix elements and controlled FT at finite $x$. This ensures that LaMET can provide a point-by-point calculation—rather than fitting—of the $x$-dependence of parton distributions, which resolves the inverse problem faced by the SDF due to limited range of accessible leading-twist correlations. In addition, we commented on the arguments and data analysis in Ref. [1], and showed that the considerable FT model uncertainty is mainly caused by the insufficient data precision in the sub-asymptotic region, which is not due to the LaMET approach itself and can be systematically improved. Moreover, the observed discrepancies among different extrapolation models are still well bounded by the theoretical estimate from the asymptotic extrapolations, thus showing that the FT is under control. This demonstrates that even when the data quality is not ideal, the asymptotic extrapolation in LaMET is still the most reliable approach for completing the FT and providing realistic error estimates. Meanwhile, re-framing the extrapolation as an inverse problem and adopting a data-driven approach with minimal physical constraint is likely to produce unnecessarily conservative errors.
@@ Acknowledgements.
We thank Daniel Hackett, Yizhuang Liu, Swagato Mukherjee, Fei Yao, and Ismail Zahed for valuable discussions. We also thank Fei Yao for sharing the data of nucleon transversity quasi-PDF. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704 and Contract No. DE-AC02-06CH11357, the Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond, and the Quark-Gluon Tomography (QGT) Topical Collaboration with Award DE-SC0023646; the National Science and Technology Council, Taiwan, under Grants No. 113-2112-M-002-012; This work is supported in part by Natural Science Foundation of China under grant No. 12125503, 12335003, 12375069, 12205106. QAZ is also supported by the Fundamental Research Funds for the Central Universities. JH is supported by Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008. YS is also supported by the 2024-2025 JSA/Jefferson Lab graduate fellowship.

### v2

@@ IV.4 Asymptotic Analysis Provides Realistic Error Estimate Regardless of Data Quality
The lattice correlation functions include statistical errors that are correlated and propagate in the FT. Since FT is a linear operation, there are mature statistical methods to handle the error propagation [266]. In addition to statistical errors, the FT also includes model uncertainty arising from the extrapolation. Another main point advocated in Ref. [1] is that when the uncertainty from extrapolation is non-negligible, it cannot be reliably quantified. We disagree. As we show here, the uncertainty can be estimated within asymptotic extrapolation and does not constitute the IP encountered in SDF.
As explained above, a non-negligible model uncertainty is the outcome of insufficient data quality in the sub-asymptotic region, which can be reduced by improving the data precision. To quantify this uncertainty, it suffices to show that it is both bounded from above and systematically reducible. The upper bound can be estimated by evaluating the contribution of the extrapolated region to the Fourier integral. One bound has been derived in the appendix of Ref. [97] as
@@
where $|h(z,P;\lambda_{L})|_{\rm max}$ is the maximal value of $|h(z,P)|$ for $\lambda_{L}\leq\lambda<\infty$  [^5] We have slightly adapted the derivation in Ref. [97], which assumed that $|h(\lambda,P)|$ decays monotonically for $\lambda>\lambda_{L}$. In our version, the derivation also allows for oscillations in this region., and $N_{x}$ is an integer which satisfies $|h^{\prime}(\lambda_{L}+2\pi N_{x}/x)|\ll|x|$ so that the contribution to the Fourier integral from $\lambda>\lambda_{L}+2\pi N_{x}/x$ is negligible.
@@
The upper bound in Eq. (12) applies to both exponential and algebraic decays, typically requiring $N_{x}\sim\mathcal{O}(1)$ in the former and $N_{x}\gg 1$ in the latter to saturate the Fourier integral. This contrast is expected: for the same $\lambda_{L}$, the extrapolated region contributes much less to the Fourier integral in the case of exponential decay, leading to a correspondingly smaller model uncertainty. For exponential decays, this upper bound is usually comparable to or smaller than the other uncertainties in the LaMET analyses, making it a reasonable estimate of the FT model uncertainty. Moreover, it excessively estimates the true uncertainty, as it assumes that $|h(z,P)|$ can vary arbitrarily between 0 and its maximum value $|h(z,P;\lambda_{L})|_{\rm max}$ at each $z$. With a more accurate derivation of the asymptotic behavior of spatial correlations [242], a significantly tighter bound on the model uncertainty can be established for the error budget.
@@
Since the RBF extrapolations in Ref. [1] include an exponential-decay component, we assume that they satisfy the conditions used to derive Eq. (12). Therefore, we may apply the same upper bound to the RBF extrapolations as to the Exact exp extrapolation. The FT model uncertainty can be estimated by the variation of central values across different models, which should persist even when the statistical errors approach zero. Because the extrapolations using the RBF kernel do not satisfy the third physical condition in Sec. III.3, we set $|h(z,P;\lambda_{L})|_{\rm max}=0.1$, reflecting the amplitude of the matrix elements in the range $0.75\,\text{fm}<z<1.03\,\text{fm}$. At $x=0.5$, taking $N_{x}=1$ is sufficient to saturate the Fourier integral, yielding an upper bound of $\sim 0.25$ on the model uncertainty. Meanwhile, according to Figure 3 of Ref. [1], the quasi-PDFs obtained from different models are consistent within errors at finite $x$, despite that their central values differ by an amount that generally decreases with increasing $x$, which is predicted by the upper bound in Eq. (12). See Figure 5 of Ref. [1]. At $x=0.5$, the spread among central values is around 0.2, implying an uncertainty of about 0.1, which is still bounded by the estimate from Eq. (12).
@@
Moreover, with the precise data in Fig. 4(a), the quasi-PDFs from different extrapolations show significantly better agreement, as illustrated in Figs. 4(b) and 4(c). The variation around the mean of the Exact exp model is less than 0.05 at $x=0.5$, which is again bounded by the estimated error of $\sim 0.06$, obtained using $|h(z,P;\lambda_{L})|_{\rm max}=0.025$ and $N_{x}=1$. Notably, the uncertainties of the GPR fits are much larger, reflecting the uncontrolled model dependence of the GPR method, as the latter can violate the physical conditions for asymptotic extrapolation. See the discussion in Sec. IV.3.
@@ IV.5 Exponential Decay is the Key to Controlling FT Errors in LaMET
Ref. [1] also made the claim that exponential decay plays a minimal role in controlling the FT at moderate to large $x$, so that the IP in LaMET and SDF is essentially the same. This claim is based on the observation that different $z$- and $x$-space reconstruction methods yield similar results for the quasi-PDF, regardless of whether exponential decay is imposed.
However, the logic behind that conclusion is in fact reversed. It is the asymptotic decay of the pseudo data—not the extrapolation model—that ensures control of the FT. This is why similar results can still be obtained even when the decay behavior is not explicitly built into the model, exactly the first issue with the data-driven methods discussed in Sec. IV.3. To match the exponentially decaying data, the extrapolation must also satisfy the corresponding physical behavior, otherwise the systematic errors cannot be reliably estimated.
In contrast, the matrix elements in SDF are limited to $\lambda\lesssim 4.0$—far from the asymptotic region—and decay much more slowly as a power law, which inevitably leads to significantly larger variations across different extrapolation models, as shown in Fig. 2. Imposing any model, even those incorporating exponential decay, does not resolve the IP in this case.
@@ V Conclusion
In this work, we have explained why the asymptotic analysis in LaMET enables a reliable extrapolation of the long-range lattice matrix elements and controlled FT at finite $x$. This ensures that LaMET can provide a point-by-point calculation—rather than fitting—of the $x$-dependence of parton distributions, which resolves the IP faced by the SDF due to limited range of accessible leading-twist correlations. In addition, we commented on the arguments and data analysis in Ref. [1], and showed that the considerable FT model uncertainty is mainly caused by the insufficient data precision in the sub-asymptotic region, which is not due to the LaMET approach itself and can be systematically improved. Moreover, the observed discrepancies among different extrapolation models are still well bounded by the theoretical estimate from asymptotic extrapolations, thus showing that the FT is under control. This demonstrates that even when the data quality is not ideal, the asymptotic extrapolation in LaMET remains the most reliable approach for completing the FT and providing realistic error estimates. Meanwhile, re-framing the extrapolation as an IP and adopting a data-driven approach with minimal physical constraint is likely to produce unnecessarily conservative errors.
@@ Acknowledgements.
We thank Daniel Hackett, William Good, Yizhuang Liu, Swagato Mukherjee, Fei Yao, and Ismail Zahed for valuable discussions. We also thank Fei Yao for sharing the data of nucleon transversity quasi-PDF. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through Contract No. DE-SC0012704 and Contract No. DE-AC02-06CH11357, the Scientific Discovery through Advanced Computing (SciDAC) award Fundamental Nuclear Physics at the Exascale and Beyond, and the Quark-Gluon Tomography (QGT) Topical Collaboration with Award DE-SC0023646, and the National Science and Technology Council, Taiwan, under Grants No. 113-2112-M-002-012. This work is supported in part by Natural Science Foundation of China under grant No. 12125503, 12335003, 12375069, 12205106. QAZ is also supported by the Fundamental Research Funds for the Central Universities. JH is supported by Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008. YS is also supported by the 2024-2025 JSA/Jefferson Lab graduate fellowship.
Appendix A Response to arXiv:2506.24037
After the appearance of this paper on arXiv, the same authors of Ref. [1] published a comment in arXiv:2506.24037 [240], where they clarified that they “did not conclude in Ref. [1] that it is not possible to quantify or control the uncertainties in the LaMET inverse problem”. Taken literally, this clarification acknowledges that the FT uncertainty can indeed be controlled and quantified in LaMET. To us, this is the central issue that distinguishes LaMET from SDF, since the latter does not have control over the uncertainty in $x$-space with the same contemporary lattice setup. With this consensus, the quantification of the FT uncertainty in LaMET becomes a practical issue.
However, the authors of Ref. [1] adopted a definition of the IP broad enough to encompass LaMET as well as most standard lattice analyses, even though these are not commonly regarded as such by the lattice community. After all, the categorical statements made in Ref. [240] are that GPR is better than the asymptotic extrapolation in quantifying the FT uncertainty in LaMET, and that exponential decay has a negligible impact on the GPR reconstruction of the $x$-dependence. Moreover, the authors promoted GPR as a reliable approach for SDF as well, which underscores their statement in Ref. [1] that LaMET and SDF share a similar IP, contrary to the main point of this paper.
First of all, as explained in Sec. II, the uncertainty in SDF cannot be controlled due to the intrinsic limitation of effective data with the contemporary lattice setup. Even if the GPR reconstruction yields results consistent with the true value in LaMET, this does not imply that the method is similarly reliable in SDF.
Second, we clarify that our proposal is not merely to perform an extrapolation using a rigid model, as claimed in Ref. [240], but to also provide an estimate of the associated uncertainty. While it is true that not all LaMET analyses in the literature have done so, in many cases this is because the uncertainty is either negligible or not dominant compared to other sources of systematics. Nevertheless, as we argued in Sec. III.3, a reliable estimate of the FT uncertainty can be obtained by incorporating knowledge of the asymptotic behavior of the correlation function. Key quantities such as the decay constant and the extent of the sub-asymptotic region can be extracted from high-precision zero-momentum matrix elements, which are frame-independent and can serve as priors at nonzero momentum. This approach removes the arbitrariness in implementing the first physical condition, which was criticized in Ref. [240]. Furthermore, since the correlation function admits a controlled asymptotic expansion with subleading terms that decay more rapidly, one can estimate the uncertainty from the missing contributions in the same spirit as in EFT calculations [242].
Third, the GPR reconstruction, at least in its current form in Refs. [224, 1, 240], does not allow for a reliable quantification of the FT uncertainty in LaMET. Before elaborating on why, we note that a meaningful uncertainty quantification entails an upper bound that can be systematically improved. This upper bound can be rigorously derived from the asymptotic expansion of the correlation functions  [^6] The conservative estimate in Eq. (12) is obtained by encoding only the leading exponential decay while leaving the remaining $z$-dependence unspecified, so $N_{x}$ remains indefinite, though it is expected to be close to 1. However, with more precise knowledge of the asymptotic behavior [242], this upper bound can be tightened without arbitrariness., which was acknowledged in Ref. [240] as the authors “have never doubted that one can produce a reasonable estimate on the maximal uncertainty of the $x$-reconstruction”. Meanwhile, GPR uses smoothness of the quasi-PDF as the sole physical regulator for solving the FT problem. Unlike the asymptotic decay of correlation functions, this notion of “smoothness” is vague and lacks a clear foundation in quantum field theory. In practice, it is primarily governed by the hyperparameters that must be chosen on a case-by-case basis [240], yet the associated model uncertainty still needs to be estimated. Most importantly, smoothness alone does not enable a rigorous derivation of an upper bound on this uncertainty, so its reliable quantification cannot be made.
Finally, the authors continued to use reversed logic in Ref. [240] in discussing the role of exponential decay in the extrapolation. It is the exponential decay of the matrix elements themselves, not the reconstruction method, that ensures the FT uncertainty admits a systematically improvable upper bound and is therefore under control  [^7] Note that exponential decay is a sufficient but not necessary property for the FT to be under control, as other rapidly decaying features would also suffice. However, in LaMET, it is the only relevant physical property of the correlation function.. The small difference observed between GPR models with and without exponential decay is a consequence of this property, which relaxes the model space for extrapolation. A similar observation was also made in Ref. [97], where both exponential and power-law decay models produced consistent results. However, when it comes to quantifying the FT uncertainty, a reliable estimate is only possible with knowledge of the asymptotic behavior of the correlation functions.
The authors of Ref. [240] also cited the deep-inelastic scattering structure function as a counterexample to our statement that the FT in LaMET is an FP. By expressing the PDF as a convolution of the matching coefficient with the structure function—i.e., as an FP—they argued that this contradicts the paradigm that PDF fitting is an IP. However, the key difference lies in the premise: so far one has not obtained the structure functions at a fixed $Q^{2}$ over a sufficiently wide range of the Bjorken variable $x_{B}$ for a reliable interpolation. In contrast, the lattice correlation functions can be computed at fixed $P^{z}$ within a wide range of $z$ and reliably extrapolated to $z\to\infty$.
Regarding “point-by-point” calculation of the PDF in LaMET, we are not referring to point-like features. Rather, our emphasis is on LaMET’s capability to calculate, rather than fit, the PDFs. As explained in Sec. II, any sharp features of the quasi-PDF or PDF must either be detected or lie within the overall uncertainty band; otherwise, their presence would violate the EFT power expansion or asymptotic decay of the correlation functions. In the example shown in Figure 4 of Ref. [240], the hypothetical bump at $x=0.5$ corresponds to an oscillatory mode that decays too slowly in coordinate space, which can be ruled out by physical constraints.